# Reinforcement Learning with Dynamic Multi-Reward Weighting for Multi-Style Controllable Generation

### v1

@@ Abstract
Style is an integral component of text that expresses a diverse set of information, including interpersonal dynamics (e.g. formality) and the author’s emotions or attitudes (e.g. disgust). Humans often employ multiple styles simultaneously. An open question is how large language models can be explicitly controlled so that they weave together target styles when generating text: for example, to produce text that is both negative and non-toxic. Previous work investigates the controlled generation of a single style, or else controlled generation of a style and other attributes. In this paper, we expand this into controlling multiple styles simultaneously. Specifically, we investigate various formulations of multiple style rewards for a reinforcement learning (RL) approach to controlled multi-style generation. These reward formulations include calibrated outputs from discriminators and dynamic weighting by discriminator gradient magnitudes. We find that dynamic weighting generally outperforms static weighting approaches, and we explore its effectiveness in 2- and 3–style control, even compared to strong baselines like plug-and-play model. All code and data for RL pipelines with multiple style attributes will be publicly available. [^1] https://github.com/minnesotanlp/multi-style-control
Reinforcement Learning with Dynamic Multi-Reward Weighting
for Multi-Style Controllable Generation
@@ 1 Introduction
Textual style is an important component of communication that conveys information not included in the text’s raw semantic content (Hovy, 1995). Consequently, it is vital that large language models are well-equipped to understand and apply styles themselves.
@@
Prior work has explored the domain of controlled generation, in which the goal is for a generative language model to continue the generation of a given prompt with text of a specified style [^2] Style transfer, on the other hand, involves paraphrasing given text to a target style without altering its original meaning. However in practice, text frequently contains not only a single style, but a combination of styles. For example, consider being asked to give feedback to a colleague at work: both formal and positive styles would be appropriate. On the other hand, if speaking with a friend about a movie, both informal and positive styles are likely to be useful. Examples of generations following multiple styles models can be found in Figure 1.
@@
Especially as large language models (LLMs) grow in capability and popularity, it is desirable to include fine-grained control of the styles included in the model’s output. For instance, in almost all cases, toxicity and hate speech must be tightly controlled such that the model does not produce harmful output. At the same time, in response to the user’s preferences or the application, it can be beneficial for the LLM to simultaneously control additional attributes such as humor, formality, or the use of figurative language. In order to achieve these goals reliably, techniques for robust multi-style control are needed. [^3] While some sociolinguistics theories distinguish between textual style and textual attributes, in this work, we follow the common convention in recent NLP papers of broadly using ‘style’ to encompass both of these ideas Jin et al. (2022).
Controlling for more than one style during generation is an under-investigated area, with prior work focusing on controlling for a single style, or a style and a target topic(s) (Keskar et al., 2019; Liu et al., 2022). Our goal in this work is towards a language model controlled for multiple styles, where each style is specified by a discriminator. The appeal of achieving multi-style control by leveraging single style discriminators is in its simplicity: pre-trained single style discriminators are readily available, and otherwise relatively easily to train. In contrast, approaches that would require obtaining or annotating large amounts of text for multiple styles are time- and labor-intensive.
An increasingly common approach to satisfying multiple desiderata for generations is reinforcement learning (RL), with each individual desideratum contributing to the reward. In this work, we explore this approach for multi-style controlled generation. As textual styles are often not orthogonal to one another, combining multiple styles into a single reward formulation is especially difficult. Styles also interact with each other in complex ways, and certain combinations – for instance, disgust and positivity – occur very rarely and are difficult to combine (Kang and Hovy, 2021). Finally, annotated data for many styles is scarce, and multi-style annotations for texts are extremely limited.
@@
We implement several strategies for formulating this multi-style reward. Our evaluations indicate that dynamically weighting each component outperforms static weighting, and we also find that simple steps such as confidence calibration and binarization of style discriminator output can improve model performance. We also implement a custom plug-and play pipeline (Dathathri et al., 2020) for comparison. In inspecting the generations of our two- and three-style controlled models, we also observe that style combinations that occur more frequently in human text appear to be easier for models to learn to generate.
@@
This is to our knowledge a first-of-its-kind work for controlling multiple styles with RL. Work on multi-reward formulation for RL is especially relevant given modern alignment techniques that incorporate multiple types of human feedback, as recent work investigates composing multiple types of human feedback – e.g. helpfulness, correctness – into the reward function Wu et al. (2023); Rame et al. (2023). Our approach is extensible to this line of work as well.
Specifically, our contributions are:
  • A first-of-its-kind comparison of the efficacy of several reward formulations for RL-based multi-style control.
  • A novel dynamic weighting formulation that shows strong control while maintaining generation quality.
  • An examination of the side-effects of style control, outlining scenarios in which factuality suffers and un-controlled styles shift.
@@ 2 Related Work
Controlled-Text Generation
Several works have investigated approaches to controlling the generations of language models with respect to various attributes. CTRL Keskar et al. (2019) retrains a class-conditional language model conditioned on many control codes to guide generations. A more lightweight approach, PPLM Dathathri et al. (2020) does not fine-tune the language model but instead uses gradients from attribute classifier(s) to guide the language model’s hidden state towards the desired state during generation. Other approaches that learn a discriminator to guide the generation of some pre-trained language model include GeDI, which computes control codes and anti-control codes for all possible next tokens Krause et al. (2020), or Attribute Alignment, which learns an alignment function Yu et al. (2021) infuse attribute representations into a pre-trained language model to guide generation.
@@
Figure 2: Three different approaches to combine multiple rewards effectively in RL training. $d_{i}(x)$ refers to the attribute from discriminator $i$ on input text $x$. We find dynamic weighting (our proposal) is most effective for multi-style control.
Our work is most related to Dathathri et al. (2020) and Liu et al. (2022), which learns some external discriminator to distinguish whether the generated text holds the desired attributes to retrain a contrastive generator (CTRL). In contrast, our work learns a set of discriminators that are combined via some reward formulation to fine-tune our model via reinforcement learning to control generation for the desired styles simultaneously.
Other works have proposed controlling the generation of language models via prefix-tuning (Li and Liang, 2021; Qian et al., 2022) by freezing model parameters and only training a continuous set of task- or style-specific "prefixes" that are prepended to steer generation.
Reinforcement Learning
Fine-tuning language models via reinforcement learning is another approach to controlling for specific attributes. Gong et al. (2019) use a style classifier model to provide a target style reward (style, content preservation, and fluency) for text style transfer. Subsequently, Upadhyay et al. (2022) build on this by considering token-level dense rewards and taking the weighted sum of these rewards that were heuristically determined to update the policy. Similarly, Luo et al. (2019) propose a dual reinforcement learning framework to learn two rewards (style accuracy and content preservation) by taking the harmonic mean of the rewards. However, these works focus on the textual style transfer task, whereas our work focuses on multi-style controllable generation.
Learning from preference data
@@
More recently, various approaches to fine-tuning language models via preferences Ziegler et al. (2020) have seen success in guiding text generations to be more aligned with desired attributes. Notably, Ramé et al. (2023) and Wu et al. (2023) train an initial model via supervised fine-tuning on a target dataset and then train separate reward models from human preference data that guide models during the RL stage. Other works have explored aligning models towards specific attributes by modeling the reward function as a preference model Rafailov et al. (2023) to bypass the need for explicitly calculating a reward function, reducing the task to a maximum likelihood objective.
@@ 3 Proposed Method: Multi-reward Control in RL
Figure 3: Pipeline for three-style control using fine-tuning Llama 7B model with PPO. We investigate several techniques for integrating feedback from style discriminators into the reward.
@@ 3.1 Preliminary Formulation: Style Control using Reinforcement Learning
Reinforcement learning for language models frames the generative language model as a policy network, $\pi_{\theta}$. The policy is a probability distribution over all possible actions (i.e., vocabulary tokens) that determines $a_{t}$, i.e. the action to be taken at timestep $t$ given the state $x_{t-1}$. We can then generate tokens based on $\pi_{\theta}(a_{t}|x_{t-1})$. Reinforcement learning also introduces a reward function $R$ that takes in a state $x_{t}$ and outputs a scalar valuation of that state. (Reinforcement learning for language models generally uses sparse rewards, i.e. a reward is only computed once the full sequence of tokens have been generated.) The objective of reinforcement learning, then, is to learn a new policy $\pi_{\theta^{\prime}}$ such that the policy maximizes the expected value of $R$. Controlling a single style using RL fine-tuning typically formulates the reward using a discriminator $d$ for the target style, i.e. the reward is $R(x)=d(x)$.
@@ 3.2 Multi-Reward Control Formulations
When constructing a reward formulation that combines outputs from multiple style discriminators, we consider two important aspects. The first is the discriminator output itself: we want to find a way to express the output, e.g. via some transformation, that provides a strong and consistent reward signal.
The second aspect we consider is the how to weigh the rewards when combining them. This is important because training can become intractable if the signal from one discriminator dominates. Consider, for example, a 2-style softmax reward formulation. Imagine that our starting policy is such that there is a near-zero likelihood of producing a generation that results in high softmax scores for both $d_{1}$ and $d_{2}$, but there is a high likelihood of producing generations that results in a high score for $d_{1}$. The policy will quickly discover how to move in a direction that maximizes scores from $d_{1}$, and move away from exploring the states that result in more balanced rewards. Therefore, we want to combine the discriminators in a way that encourages balancing results from all discriminators.
Motivated by these considerations, we explore multiple approaches to calculate a reward $R$ for a generation $x$ by combining output from the attribute discriminators $d_{1},d_{2},...d_{n}$ with target styles $k_{1},...,k_{n}$, and we write the logit value for the target class as $d_{i}(x)_{k_{i}}$. We consider the following multi-reward formulations:
@@
We take the logits of the target class from the discriminator output, $R=\sum_{i=0}^{n}d_{i}(x)_{k_{i}}$.
@@
We take the softmax $\sigma$ values of the target class from the discriminator output, $R=\sum_{i=0}^{n}\sigma(d_{i}(x))_{k_{i}}$.
@@
When discriminators are unsure of their prediction, the reward signal becomes noisier, potentially hampering the policy learning process. Thus, we consider binarized rewards to make the signals much more discrete and emphasized.
$R=\sum_{i=0}^{n}w_{i},\quad w_{i}=\begin{cases}1&\sigma(d_{i}(x))_{k_{i}}\geq 0.5\\ -1&\text{ otherwise }\end{cases}$ (1)
@@
$R=\sum_{i=0}^{n}\frac{d_{i}(x)_{k_{i}}}{\texttt{temperature}_{i}}$ (2)
@@
Labels
positive,          negative
formal,         informal
ironic,         not ironic
Fear, anger, joy, sadness, disgust, surprise, neutral
toxic,          non-toxic
Dataset
SST2
GYAFC
SemEval
go-emotions
Jigsaw
F1 (Macro)
0.90
0.90
0.78
0.53
0.80
@@
$R=\sum_{i=0}^{n}w(1-\sigma(d_{i}(x))_{k_{i}}$ (3)
$w=\begin{cases}-\mathtt{grad\_norm}_{i}&\text{ if }\sigma(d_{i}(x))_{k_{i}}>0.5\\ \mathtt{grad\_norm}_{i}&\text{ otherwise }\end{cases}$ (4)
@@ 4.1 Base Models
We use LLaMA2 7B Touvron et al. (2023) as the base model for both the discriminators and the RL pipeline as shown in Figure 3. The discriminators are a classification head consisting of a fully connected linear layer. We train discriminators for sentiment, formality, irony, emotion, and toxicity using the SST2, GYAFC, SemEval-Irony, Go-emotions, and Jigsaw Toxicity datasets Socher et al. (2013); Rao and Tetreault (2018); Van Hee et al. (2018); Demszky et al. (2020); Jigsaw . For details see Table 1.
We train these custom discriminators rather than using existing classifiers because classifiers with the same architecture are needed for the PPLM Dathathri et al. (2020) pipeline. We then use these same custom discriminators in the RL fine-tuning for better comparison with PPLM. The custom discriminators are evaluated on the test sets with macro F1 and achieve results comparable to those published in the original dataset papers.
4.2 Configurations of RL Training
Figure 3 illustrates the pipeline for RL fine-tuning of LLaMA2. The language model generates a completion given a prompt from our dataset. The final hidden state from the model is pooled and passed to each target style’s discriminator. Scores from each discriminator are combined into a reward. We use the TRL library Von Werra et al. (2022) implementation of the proximal policy optimization (PPO) algorithm Schulman et al. (2017). PPO, given a batch of rewards and generations, updates the language model with the goal of maximizing its expected future rewards. Due to computational constraints, we also use the parameter efficient fine-tuning (PEFT, Mangrulkar et al. (2022)) library’s implementation of low rank adapters (LoRA, Hu et al. (2022)).
The PPO objective includes a penalty term for the Kullback–Leibler (KL) divergence between the fine-tuned language model and the original language model. During RL fine-tuning, we adjust KL-related hyperparameters, as different reward formulations perform better with different KL controls. We use an adaptive KL control setting and vary the initial KL coefficient in the range of $[0.2,0.4]$, picking the parameter with the best reward performance, provided that the KL divergence with the reference model does not rise above 20.

### v3

# Dynamic Multi-Reward Weighting for Multi-Style Controllable Generation

@@ Abstract
Textual style expresses a diverse set of information, including interpersonal dynamics (e.g., formality) and the author’s emotions or attitudes (e.g., disgust). An open question is how language models can be explicitly controlled so that they weave together target styles when generating text: for example, to produce text that is both negative and non-toxic. One approach to such controlled generation is multi-objective reinforcement learning (RL), but how to best combine multiple objectives in a reward function is an open question. In this paper, we investigate various formulations of multi-style rewards, including calibrated outputs from discriminators and dynamic weighting by discriminator gradient magnitudes. We find that our proposed dynamic weighting outperforms static weighting approaches with respect style control while maintaining linguistic quality, and we explore its effectiveness in 2- and 3–style control. All code and data for the RL pipelines will be publicly available. [^1] https://github.com/minnesotanlp/dynamic-multi-reward-weighting
Dynamic Multi-Reward Weighting for Multi-Style Controllable Generation
@@ 1 Introduction
Textual style is an important component of communication that conveys information not included in the text’s raw semantic content (Hovy, 1995). Consequently, it is vital that language models can understand and apply styles themselves.
@@
Prior work has explored the domain of controlled style generation, a task in which a generative language model aims to generate text with a specified style [^2] Style transfer, on the other hand, involves paraphrasing given text to a target style without altering its original meaning. However in practice, text frequently contains not only a single style, but a combination of styles Kang and Hovy (2021). For example, consider being asked to give feedback to a colleague at work: both formal and positive styles would be appropriate. On the other hand, if speaking with a friend about a movie, both informal and positive styles are likely to be useful. Examples of generations following multiple styles models can be found in Figure 1.
@@
Especially as large language models (LLMs) grow in capability and popularity, it is desirable to include fine-grained control of the styles in LLM outputs. For instance, in almost all cases, toxicity and hate speech must be tightly controlled such that the model does not produce harmful output. At the same time, in response to the user’s preferences or the application, it can be beneficial for the LLM to simultaneously control additional attributes such as humor, formality, or the use of figurative language. In order to achieve these goals reliably, techniques for robust multi-style control are needed. [^3] While some sociolinguistics theories distinguish between textual style and textual attributes, in this work, we follow the common convention in recent NLP papers of broadly using ‘style’ to encompass both of these ideas Jin et al. (2022).
Controlling for more than one style during generation is an under-investigated area, with prior work focusing on controlling for a single style, or a style and a target topic(s) (Keskar et al., 2019; Liu et al., 2022). In this work, we investigate the use of Reinforcement learning (RL) for controlling multiple styles. RL approaches satisfy multiple desiderata for generations by employing a reward function in which each individual desideratum contributes to the reward; this approach is recently gaining more interest in the alignment literature (see e.g., fine-grained reinforcement learning from human feedback (RLHF) from Wu et al. (2023)). In this work, we apply a similar approach for multi-style controlled generation, in which “style scores” from individual style discriminators are combined into a single reward function during RL fine-tuning.
@@
The optimal approach to combining multiple reward signals into a reward function is an open question. To further explore this question, we implement several strategies for formulating the multi-style reward, including a novel dynamic-weighting approach. Interestingly, our evaluations indicate that dynamically weighting each component outperforms static weighting, and we also find that simple steps such as confidence calibration and binarization of style discriminator output can improve model performance. We also implement a custom plug-and play pipeline (Dathathri et al., 2020) for comparison.
@@
This is to our knowledge a first-of-its-kind work investigating multi-style controllable generation through an RL lens with a new reward shaping approach via dynamic gradient-based weighting. Work on multi-reward formulation for RL is especially relevant in the current landscape, given the modern alignment techniques that incorporate multiple axes of human feedback, as recent work investigates composing multiple types of human feedback – e.g., helpfulness, correctness – into the reward function Rame et al. (2023).
@@ 2 Related Work
Figure 2: Three different approaches to combine multiple rewards effectively for reinforcement learning. $d_{i}(x)$ refers to the attribute from discriminator $i$ on input text $x$. We find dynamic weighting (our proposal) is particularly effective for multi-style control.
Controlled Text Generation
Methods for controlled generation can generally be grouped into three main categories: fine-tuning, retraining, and post-processing Zhang et al. (2023). Post-processing approaches are the most lightweight and involve applying transformations during decoding, rather than making any adjustments to model weights themselves. Examples of such methods include plug and play, or PPLM Dathathri et al. (2020)), which uses gradients from an attribute classifier to guide the language model’s hidden state; generative discriminators (GeDI) which compute control codes and anti-control codes for all possible next tokens Krause et al. (2020); and Attribute Alignment, which learns an alignment function Yu et al. (2021) infuse attribute representations into a pre-trained language model to guide generation. Prefix-tuning (Li and Liang, 2021; Qian et al., 2022) can also guide generation by prepending task- or style-specific “prefix” vectors. Retraining (or refactoring) methods involve retraining language models from the ground up on the control task; for example, CTRL Keskar et al. (2019) retrains a class-conditional language model conditioned on many control codes to guide generations. Another retraining approach is Cev-LM Moorjani et al. (2024), a prototype-then-edit semi-autoregressive language model that applies edit vectors in the latent space.
@@
Our work falls under the fine-tuning category. Fine-tuning methods adjust parameters of a pre-trained LLM toward fulfilling the desired controls. Reinforcement learning (RL) is a common fine-tuning approach for controlled text generation Zhang et al. (2023), e.g., Gong et al. (2019) use a style classifier model to provide a target style reward, and Upadhyay et al. (2022) use token-level dense rewards and taking the weighted sum of these rewards that were heuristically determined to update the policy. Other works align models towards specific attributes by modeling the reward function as a preference model Rafailov et al. (2023) to bypass the need for explicitly calculating a reward function, reducing the task to a maximum likelihood objective. More recently, various approaches to fine-tuning language models via human preferences Ziegler et al. (2020); Ouyang et al. (2022) have seen success in guiding text generations to be more aligned with desired attributes.
Finally, we point out that textual style transfer is related to controlled generation, but it is a distinct task that involves transforming an input text’s style while preserving the semantics. Recent work in style transfer include Steering Vectors (Subramani et al., 2022), which inject steering vectors into the model during decoding. Variations such as ActAdd Turner et al. (2023) and activation engineering Konen et al. (2024) have also been proposed.
Multi-Objective Rewards
Recent work explores how to incorporate reward signals from multiple sources into reward functions, particularly in alignment literature involving RLHF. Notably, recent work investigates multi-objective RLHF by training separate reward models from human preference data and linearly combining those rewards Wu et al. (2023); Rame et al. (2024); Ramé et al. (2024). In this work, we apply a similar approach for controlled style generation, and we also experiment with additional approaches to reward combination.
Inverse RL
Our work in dynamically shaping the weights of the reward function is also adjacent to Inverse Reinforcement Learning (IRL), which learns both a policy and a reward function. IRL alternates policy update steps with reward model update steps, requiring expert demonstrations for the reward update step. For example, Ghosh et al. (2021) use IRL for Table-to-Text generation, and their reward function takes several table descriptors as inputs; and Fu et al. (2023) use IRL for the text summarization task, with a reward function that combines various sub-rewards like saliency and coverage. However, we clarify a distinction between our work and IRL: whereas IRL explicitly learns a reward function, our dynamic weighting method is best characterized as reward shaping for online policy update. Specifically, we do not use expert demonstrations to optimize the weights of the reward function, but instead dynamically "shape" the coefficients of the fine-grained reward function components during the policy update step by inspecting the gradients for each component.
@@ 3 Proposed Method: Multi-Reward Control in RL
Figure 3: Pipeline for three-style control using fine-tuning Llama 7B model with Proximal Policy Optimization (PPO). We investigate several techniques for integrating feedback from style discriminators into the reward.
@@ 3.1 Preliminary Formulation: Style Control using Reinforcement Learning
Reinforcement learning for language models frames the generative language model as a policy network, $\pi_{\theta}$. The policy is a probability distribution over all possible actions (i.e., vocabulary tokens) that determines $a_{t}$, i.e., the action to be taken at timestep $t$ given the state $x_{t-1}$. We can then generate tokens based on $\pi_{\theta}(a_{t}|x_{t-1})$. Reinforcement learning also introduces a reward function $R$ that takes in a state $x_{t}$ and outputs a scalar valuation of that state. (Reinforcement learning for language models generally uses sparse rewards, i.e., a reward is only computed once the full sequence of tokens have been generated.) The objective of reinforcement learning, then, is to learn a new policy $\pi_{\theta^{\prime}}$ such that the policy maximizes the expected value of $R$. Controlling a single style using RL fine-tuning typically formulates the reward using a discriminator $d$ for the target style, as in $R(x)=d(x)$.
@@ 3.2 Multi-Reward Control Formulations
When constructing a reward formulation that combines outputs from multiple style discriminators, we consider two important aspects. The first is the discriminator output itself: we want to find a way to express the output, e.g., via some transformation, that provides a strong and consistent reward signal.
The second aspect we consider is the how to weigh the combined reward. This is important because training can become intractable if the signal from one discriminator dominates. For instance, consider a 2-style softmax reward formulation, and imagine our starting policy is such that there is a near-zero likelihood of producing a generation that results in high softmax scores for both $d_{1}$ and $d_{2}$, but there is a high likelihood of producing generations that results in a high score for $d_{1}$. The policy will quickly discover how to move in a direction that maximizes scores from $d_{1}$ and move away from exploring the states that result in more balanced rewards. Instead, we want to combine the discriminators in a way that encourages balancing all outputs.
Motivated by these considerations, we explore multiple approaches to calculate a reward $R$ for a generation $x$ by combining output from the attribute discriminators $d_{1},d_{2},...,d_{n}$ with corresponding target styles $k_{1},k_{2},...,k_{n}$, and we write the logit value for the target class as $d_{i}(x)_{k_{i}}$. We consider the following multi-reward formulations:
@@
We take the logits of the target class from the discriminator output, $R=\sum_{i=1}^{n}d_{i}(x)_{k_{i}}$.
@@
We take the softmax $\sigma$ values of the target class from the discriminator output, $R=\sum_{i=1}^{n}\sigma(d_{i}(x))_{k_{i}}$.
@@
When discriminators are unsure of their prediction, the reward signal becomes noisier, potentially hampering the policy learning process. Thus, we consider binarized rewards to make the signals more discrete and emphasized.
$R=\sum_{i=1}^{n}w_{i},\quad w_{i}=\begin{cases}1&\sigma(d_{i}(x))_{k_{i}}\geq 0.5\\ -1&\text{ otherwise }\end{cases}$ (1)
@@
$R=\sum_{i=1}^{n}\frac{d_{i}(x)_{k_{i}}}{\texttt{temperature}_{i}}$ (2)
@@
Labels positive,          negative formal,         informal ironic,         not ironic Fear, anger, joy, sadness, disgust, surprise, neutral toxic,          non-toxic
Dataset SST2 GYAFC SemEval go-emotions Jigsaw
F1 (Macro) 0.90 0.90 0.78 0.53 0.80
@@
$R=\sum_{i=1}^{n}w(1-\sigma(d_{i}(x))_{k_{i}})$ (3)
$w=\begin{cases}\mathtt{grad\_norm}_{i}&\text{ if }\sigma(d_{i}(x))_{k_{i}}>0.5\\ -\mathtt{grad\_norm}_{i}&\text{ otherwise }\end{cases}$ (4)
@@
We compose all reward shaping methods by taking the convex combination of each discriminator output, with only the gradient scaling method differing. For example, the binarized, confidence, and calibration methods are all follow the formulation with given $d_{1}$ and $d_{2}$ (where each $d_{n}$ refers to some style classifier)
$R_{\text{final}}=\alpha\cdot d_{1}(x)+(1-\alpha)\cdot d_{2}(x)$ (6)
where each $d_{1}$ and $d_{2}$ rewards are shaped via each respective method. Here we take the $\alpha$ as $1/n$ where $n$ is the number of discriminators and can be extended to any number $n$ such that $\sum_{i=1}^{n}\alpha_{i}=1$. For the gradient-weight scaling, we replace $\alpha$ with grad_normi (extendable to any $n$ number of discriminators)
$R_{\text{final}}=\texttt{grad\_norm}_{1}\cdot d_{1}(x)+\texttt{grad\_norm}_{2}\cdot d_{2}(x)$ (7)
@@ 4.1 Base Models
We use LLaMA2 7B Touvron et al. (2023) as the base model for both the discriminators and the RL pipeline as shown in Figure 3. We train discriminators for sentiment, formality, irony, emotion, and toxicity using the SST2, GYAFC, SemEval-Irony, Go-emotions, and Jigsaw Toxicity datasets Socher et al. (2013); Rao and Tetreault (2018); Van Hee et al. (2018); Demszky et al. (2020); Jigsaw . We train these custom discriminators rather than using existing classifiers because classifiers with the same base model architecture are needed for the PPLM Dathathri et al. (2020) pipeline. For better comparison with PPLM, we use these same custom discriminators in the RL fine-tuning as well. The discriminators are evaluated on the test sets with macro F1 and achieve results comparable to those published in the original dataset papers (Table 1).
4.2 Configurations of RL Fine-Tuning
Figure 3 illustrates our pipeline for RL fine-tuning. The language model generates a completion given a prompt from our dataset. The pooled final hidden state from the model is input to each target style discriminator, whose outputs are then combined into a reward. We use the TRL library Von Werra et al. (2022) implementation of the proximal policy optimization (PPO) algorithm Schulman et al. (2017). Due to computational constraints, we use low rank adapters (LoRA, Hu et al. (2022)) implemented by the parameter efficient fine-tuning (PEFT, Mangrulkar et al. (2022)) library.
The PPO objective includes a penalty term for the Kullback–Leibler (KL) divergence between the fine-tuned and original language model. During training, we use an adaptive KL control setting and do a parameter search for the initial KL coefficient in the range of $[0.2,0.4]$, eliminating any runs that result in final KL divergences over 20. (Larger KL divergence values in our experience were associated with repetitive reward hacking behaviors.)